Status Report

SOHO Daily Meeting Minutes for Friday, 11 January 2002

By SpaceRef Editor
January 11, 2002
Filed under , ,



* See the weekly meeting minutes for details on the maneuver next



Spacecraft Status: Nominal

Spacecraft Anomalies:

011/1102 CEPAC ESU Data Request Error, recovered autonomously after 4 minutes.

Accomplished Activities: VIRGO, Nominal HGA Table

Planned Activities: VIRGO, SWAN, RSL, New control star, mag = 6.5

Upcoming Operations: 01/15 SK/MM Maneuvers

01/16 MDI Continuous starts

Ground Anomalies:

  010/2110      Lost uplink lock on spacecraft receiver 2, cause
unknown. Lost command ability for three minutes.
011/0800 Swapped TGCs at D66 due to an invalid clock value.



01/10 DOY 010 – 01/11 DOY 011

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:



CDS: Nominal. Continuing observations with Lick and BBSO on AR9773,
and the usual weekend studies will be run.

UVCS: Nominal. Finishing up JOP151 then beginning coronal hole
observations in the north and synoptic.

LASCO: Nominal. C2/C3 synoptics. High cadence observations begin when
EIT bakeout starts.

EIT: Nominal. 195 Angstrom CME watch, synoptics. Bakeout starts around
22 UT.

MDI: Nominal. Full disk magnetograms and dopplergrams. Doing a tuning load
in preparation for the maneuver and the continuous.

TRACE: Nominal. Flarewatch on AR 9773.

SpaceRef staff editor.